for each of the following equation , determine whether demand is elastic, inelastic, or unitary elastic at given price.
a. Q=100-4P and P = Rs 20.
b. Q= 1500-20P and P= Rs 5.
c. P=50-0.1Q and P Rs.20.
1
Expert's answer
2020-04-10T09:06:50-0400

 Edp = Q'p*P0/Q0

a) Q'p=(100-4P)'=-4, P0=20, Q0=100-80=20,  Edp = -4*20/20=-4, demand is elastic (Edp>1)

b) Q'p=(1500-20P)'=-20, P0=5, Q0=1500-100=1400,  Edp = -20*5/1400=-0.07, demand is inelastic (Edp<1)

c) Q'p=(500-P)'=-1, P0=20, Q0=500-20=480,  Edp = -1*20/480=-0.04, demand is inelastic (Edp<1)